# An LLM Agent for Functional Bug Detection in Network Protocols

### v1

@@ III-B Phase 2: Retrieval-guided Detection
Technically, the retrieval-guided detection begins by pre-processing the RFC document and segmenting it into sections and subsections based on its structural headings. These segments are considered the semantic properties that guide the detection process. For each semantic property, the detection agent scans the implementation by following a workflow similar to a finite state machine, as illustrated in Figure 5. Specifically, it first localizes relevant functions for inspection (Localization), determines whether the implementation satisfies the RFC requirement (Detection), and, if necessary, retrieves additional context (Retrieval). When a potential functional bug is found, the detection agent examines the reasoning chain that produces the reported bug via self-criticism (Validation). In the following subsections, we provide further details on each component of the detection workflow shown in Figure 5.
@@
Retrieval. To assist the LLM in identifying functional bugs or justifying correctness, the detection agent enables demand-driven retrieval of additional program constructs. This is achieved by equipping the LLM with a set of predefined tools, which it can invoke through function calls as needed. Specifically, the agent provides three types of tools. First, Query(name) retrieves the definition of a specific data structure, macro, or global variable. This is useful when the current context references symbols whose definitions are not yet available. Second, Query_Callee(call) returns the definition of the callee function invoked at the specified call site. This is used when the current function doesn’t provide enough information to determine whether the behavior matches the RFC, and the LLM suspects that the called function contains important logic. This selective strategy enables the LLM to focus on semantically meaningful functions while avoiding unnecessary expansion of trivial or unrelated calls. Third, Query_Caller(fun) retrieves all callers of the function fun, allowing the LLM to examine how the function is used, particularly whether its preconditions are satisfied by the callers. These retrieval operations are triggered on demand, guided by the LLM’s reasoning. Benefiting from model’s planning ability, the LLMs can choose the proper tools for retrieval. The retrieved results further augment the analysis context in the Detection stage for continued analysis. The LLM prompt is shown in Figure 6(c).
@@
Validation. Due to the potential lengthy detection context, the LLMs may hallucinate and report a false positive. To mitigate hallucination, we employ self-critics [24] in the detection agent. Specifically, the detection agent reviews all available information, including the RFC section, the retrieved context, and earlier reasoning steps, to make a final judgment. If the violation is confirmed, the functional bug is reported as a true positive. Otherwise, it is discarded as a false positive. It may also uncover bugs missed during previous detection. The prompt is instantiated using the template in Figure 6(d). Our results in Section IV-E show that the self-critics can reduce the false positives by 71.9% (from 64.3% to 18.1%).
@@
Given the RFC section in Figure 1(a) and the semantic indexes in Figure 3(a), the detection agent identifies the relevant function route_lost in route.c (Figure 7). Finding its context insufficient, the agent retrieves find_best_route and send_update_resend, which select candidate routes and issue reactive updates, respectively. Notably, send_update_resend resides in a different file. With this broader context, the agent observes that only feasible routes are searched (flag = 1) and no sequence number request is made in the fallback logic, violating the starvation prevention rule. Retrieval stops here, and the bug is identified as shown in Figure 1 and detailed in Figure 3(b).
@@ IV Implementation and Evaluation
RFCScan is implemented atop AutoGen [25], a multi-agent framework designed for constructing LLM applications. To extract functional specifications, we collect RFC documents in plain text format and derive informal functional properties through structured segmentation of the textual content. We employ Claude 3.5 Sonnet [26] as the underlying LLM for RFCScan, configured with a temperature of 0.0 to enforce greedy decoding. This setting reduces randomness in model responses, thereby enhancing the consistency and reproducibility of RFCScan. To facilitate the indexing and detection agents, we implement a suite of analysis tools using Tree-sitter [27], a parsing library for different programming languages. In particular, we construct call graphs by leveraging function names and the numbers of their associated parameters or arguments. The output generated by the indexing agent is stored in a JSON file, enabling incremental analysis as the protocol implementation evolves. This design avoids redundant processing by skipping the analysis of unchanged functions, files, and modules.
To assess the performance of RFCScan, we conduct experiments to address the following research questions:
  • RQ1: How effective does RFCScan identify functional bugs in real-world network protocol implementations?
  • RQ2: How does RFCScan compare to existing approaches?
  • RQ3: What are the runtime and token costs of RFCScan?
@@ IV-A Dataset
We evaluate RFCScan on six protocol implementations from two widely-used, open-source network stacks: FRRouting [28] and lwIP [29]. FRRouting (FRR) is an Internet routing protocol suite for Linux and Unix platforms, with over 3.6K stars on GitHub. It is deployed in production by major infrastructure providers such as NVIDIA and Orange, and is distributed through mainstream Linux repositories including Debian and Fedora. lwIP is a lightweight TCP/IP stack designed for resource-constrained environments, with over 1.3K stars on GitHub. It is widely integrated into embedded systems, IoT platforms, and real-time applications, and has been adopted by vendors like Intel, Xilinx, and Freescale. Overall, both FRRouting and lwIP are actively maintained and widely used in practice, making them ideal subjects for assessing RFCScan across diverse deployment contexts. For each protocol, we collect the corresponding RFC that the implementation is based on as the reference. The protocol implementation sizes range from 5.2 KLoC to 17.3 KLoC, whereas the corresponding RFC documents span between 19 and 54 pages. The detailed information is shown in Table I.
@@
IV-B RQ1: Effectiveness of RFCScan
@@
Setup and Metrics. To answer RQ1, we evaluate RFCScan upon each protocol implementation and the corresponding RFC. For each protocol, we collect the functional bugs reported by RFCScan and manually verify whether each one reflects a real violation. Then we report the numbers of true positives and false positives, and calculate the false positive rate. Since a single underlying issue may result in multiple reported bugs—e.g., when the same property is described across multiple sections of the RFC—we group related reports into unique bugs. For each unique bug, we check the project’s issue tracker and pull requests to determine whether it is a previously unreported bug (i.e., new) or a known issue (i.e., old) that has not yet been fixed in the latest version.
TABLE III: The list of unique bugs.
ID Protocol Type Description New? Status
1 BABEL Routing Request forwarding does not prioritize feasible routes. PR Merged
2 BABEL Routing Miss check unexpired unfeasible routes to avoid starvation. PR Merged
3 BABEL Config Miss source port validation before parsing. PR Merged
4 BABEL Config Missing router-id validation in babel_get_myid. PR Merged
5 BABEL Parsing known_ae() function accepts incorrect AE values. PR Merged
6 BABEL Parsing Incorrect parsing of type field in parse_request_subtlv. PR Merged
7 BABEL Parsing Handle unknown TLVs incorrectly in parse_request_subtlv. Pending
8 BABEL Parsing Miss mandatory bit handling in parse_ihu_subtlv. Pending
9 BABEL Parsing Miss Router-Id validation for RouterId TLV. PR Approved
10 BABEL Parsing Miss Router-Id Validation for Seqno Request TLV. PR Approved
11 BFD Parsing Miss check Simple Password Authentication format. Confirmed
12 BFD Parsing Miss check Keyed MD5 Authentication format. Confirmed
13 BFD Parsing Miss check Meticulous Keyed MD5 Authentication format. Confirmed
14 BFD Parsing Miss check Keyed SHA1 Authentication format. Confirmed
15 BFD Parsing Miss check Meticulous Keyed SHA1 Authentication format. Confirmed
16 BFD Demultiplex Incorrect session lookup for demultiplexing. PR Merged
17 BFD State Miss state validation for demand bit in ptm_bfd_send. Pending
18 BFD Demultiplex Missing poll Sequence for demand bit changes. Pending
19 BFD Timer Incorrect initialization of DesiredMinTxInterval. PR Approved
20 BFD Config Miss check passive mode with zero discriminator. PR Merged
21 BFD State LocalDiag not being set. PR Merged
22 BFD State Fail to maintain session state even when a session is down. Pending
23 BFD State Fail to start or stop Demand mode at any time. Pending
24 BFD Timer Miss check increased transmission interval. PR Merged
25 RIPng Routing Incorrect address selection for non-RIPng port requests. Pending
26 RIPng Routing Miss link-local route filtering before adding into RTE. Pending
27 NHRP Routing Miss hop count validation in nhrp_packet_pull(). PR Merged
28 NHRP State Non-volatile request ID storage violation. Pending
29 NHRP State No storage/checking of uniqueness flag. Pending
30 NHRP State No rejection of duplicate unique registrations. Pending
31 NHRP State Miss uniqueness conflict detection and error handling. Pending
32 NHRP Err Handle Fail to prevent sending error indication pkts to error pkts. Pending
33 NHRP Err Handle Fail to guarantee single error per packet rule. Pending
34 NHRP State Miss Purge Request handler. Pending
35 NHRP Routing Miss loop detection for Responder Address Extension. Confirmed
36 NHRP Routing Fails to drop packets with NHS’s own source addresses. Pending
37 DHCP State Miss response handling for DHCPINFORM messages. Pending
38 DHCP State DHCPREQUEST does not reuse the secs field. Pending
39 DHCP State Incorrect hooking message type in REBINDING stage. Pending
40 DHCP Timer Violation of T1/T2 timer relationship requirement. Pending
41 IGMP Timer Incorrect timer range coverage due to randomization error. Pending
42 IGMP State Miss version state tracking and handling per interface. Pending
43 IGMP Timer Incorrect IGMPv1 Max Response Time handling. Pending
44 IGMP State Missing report version selection logic. Pending
45 IGMP State Incorrectly allow IGMP join from invalid states. Pending
46 IGMP Timer Use a 500ms delay instead of the RFC-mandated 10 seconds. Pending
47 IGMP Timer Miss Version 1 router present timeout. Pending
@@
Results. As shown in Table table II, RFCScan generates a total of 83 bug reports across six protocols, of which 67 are true positives, yielding an overall precision of 81.9%. The false positive rate remained low, with most protocols under 25%, and BFD exhibiting the lowest at only 10.0%. Grouping related bug reports, we uncovered 47 unique bugs, including 40 new bugs and 7 already reported ones. The complete bug list is shown in  Table III. To date, 20 out of 47 unique bugs have been confirmed or their patches have been approved or merged by the developers. Overall, these results demonstrate that RFCScan is effective in identifying bugs across diverse protocols. To better understand the diversity of functional bugs detected by RFCScan, we categorize the 47 unique bugs and obtain seven categories, namely Parsing, Config, Routing, Err Handle, State, Timer, and Demultiplex. As shown in Table III, most bugs fall under State (15/47) and Parsing (11/47), highlighting common pitfalls in maintaining protocol state and validating input structure. However, RFCScan also discovered bugs such as Timer misconfiguration (e.g., incorrect IGMP delays in Bug #46), Demultiplexing issues (e.g., Incorrect BFD session look up in Bug #16), and Err Handle gaps (e.g., missing error-on-error prevention in Bug #32). This variety highlights the ability of RFCScan to detect diverse semantic violations specified by the RFCs.
IV-C RQ2: Baseline Comparison
Setup and Metrics. To the best of our knowledge, no prior work targets functional bug detection in network protocols based on RFCs. We compare RFCScan with GitHub Copilot using three LLMs: Claude 3.7 Sonnet (Thinking Mode), Claude 3.5 Sonnet, and GPT-4o. All baselines are accessed via Copilot’s Workspace Chat interface in Visual Studio Code, which supports repository-level queries. We prompt Copilot to identify deviations between protocol implementations and RFC documents. To ensure a fair comparison with RFCScan, which analyzes one RFC section at a time, we provide Copilot with one section at a time in a single prompt. For GitHub Copilot with Claude 3.7 Sonnet, we manually examine all generated bug reports. However, Claude 3.5 Sonnet and GPT-4o produce hundreds of candidates, making full manual verification impractical. To avoid over-reporting and minimize false positives, we instructed each LLM baseline to only report bugs with 100% confidence. To enable a fair and scalable evaluation, we adopt a sampling-based evaluation. For each protocol, we randomly sample up to twice the number of bug reports identified by RFCScan. For example, since RFCScan generates 16 bug reports for Babel, we sample 32. If the model produces fewer than that, we evaluate all of them. Each sampled case is then manually reviewed by a domain expert. We do not compare with existing network protocol bug detection tools, as they either cannot detect functional bugs (e.g., ChatAFL [9], NetLifter [17]), or target narrow scope like parsing (e.g., ParCleanse [21], ParVAL [30]).
@@ IV-C RQ2: Baseline Comparison
Results. Table IV shows the performance of GitHub Copilot under three model configurations. – With Claude 3.7 Sonnet (no sampling), it generated 215 bug reports, of which 78 were true positives (36.3% precision), uncovering 37 unique bugs (29 new, 8 old). – Claude 3.5 Sonnet achieved 19.9% precision over 538 inconsistencies and found 26 bugs. – GPT-4o returned the most bug reports (1,315) but the lowest precision (11.4%) and the fewest bugs (16). Overall, all three baselines suffer from low precision and high validation overhead. In contrast, RFCScan achieves 81.9% precision and discovers 47 bugs in total (40 new) with far less manual effort (Table II). RFCScan consistently reports the lowest false-positive rates and the highest bug counts across all protocols. Notably, RFCScan ’s use of Claude 3.5 (a non-reasoning model) still outperforms Claude 3.7 (a reasoning model).
IV-D RQ3: Efficiency of RFCScan

### v2

@@ III-B Phase 2: Retrieval-guided Detection
Technically, the retrieval-guided detection begins by pre-processing the RFC document and segmenting it into sections and subsections based on its structural headings. From each subsection, the LLMs extract mandatory semantic properties, which serve as guides for the detection process. A single subsection may yield multiple properties depending on its content. For each semantic property, the detection agent scans the implementation by following a workflow similar to a finite state machine, as illustrated in Figure 5. Specifically, it first localizes relevant functions for inspection (Localization), determines whether the implementation satisfies the RFC requirement (Detection), and, if necessary, retrieves additional context (Retrieval). When a potential functional bug is found, the detection agent examines the reasoning chain that produces the reported bug via self-criticism (Validation). In the following subsections, we provide further details on each component of the detection workflow shown in Figure 5.
@@
Retrieval. To assist the LLM in identifying functional bugs or justifying correctness, the detection agent enables demand-driven retrieval of additional program constructs. This is achieved by equipping the LLM with a set of predefined tools, which it can invoke through function calls as needed. Specifically, the agent provides three types of tools. First, Query(name) retrieves the definition of a data structure and a macro. This is useful when the current context references symbols whose definitions are not yet available. Second, Query_Callee(call) returns the definition of the callee function invoked at the specified call site. This is used when the current function doesn’t provide enough information to determine whether the behavior matches the RFC, and the LLM suspects that the called function contains important logic. This selective strategy enables the LLM to focus on semantically meaningful functions while avoiding unnecessary expansion of trivial or unrelated calls. Third, Query_Caller(fun) retrieves all callers of the function fun, allowing the LLM to examine how the function is used, particularly whether its preconditions are satisfied by the callers. These retrieval operations are triggered on demand, guided by the LLM’s reasoning. Benefiting from model’s planning ability, the LLMs can choose the proper tools for retrieval. The retrieved results further augment the analysis context in the Detection stage for continued analysis. The LLM prompt is shown in Figure 6(c).
@@
Validation. Due to the potential lengthy detection context, the LLMs may hallucinate and report a false positive. To mitigate hallucination, we employ self-critics [33] in the detection agent. Specifically, the detection agent reviews all available information, including the RFC section, the retrieved context, and earlier reasoning steps, to make a final judgment. If the violation is confirmed, the functional bug is reported as a true positive. Otherwise, it is discarded as a false positive. It may also uncover bugs missed during previous detection. The prompt is instantiated using the template in Figure 6(d). Our results in Section IV-E show that the self-critics can reduce the false positives by 71.7% (from 63.9% to 18.1%).
@@
Given the RFC section in Figure 1(a) and the semantic indexes in Figure 3(a), the detection agent identifies the relevant function route_lost in route.c (Figure 7). Finding its context insufficient, the agent retrieves find_best_route and send_update_resend, which select candidate routes and issue reactive updates, respectively. Notably, send_update_resend resides in a different file. With this broader context, the agent observes that only feasible routes are searched (flag = 1) and no sequence number request is made in the fallback logic, violating the starvation prevention rule. Retrieval stops here, and the bug is finally identified as shown in Figure 1. The whole process is detailed in Figure 3(b).
@@ IV Implementation and Evaluation
RFCAudit is implemented atop AutoGen [34], a multi-agent framework designed for building LLM applications. To extract functional specifications, we collect RFC documents in plain text format and derive informal functional properties through structured segmentation of the textual content. We employ Claude 3.5 Sonnet [35] as the underlying LLM for RFCAudit, configured with a temperature of 0.0 to enforce greedy decoding. This setting reduces randomness in model responses, thereby enhancing the consistency and reproducibility of RFCAudit. To facilitate the indexing and detection agents, we implement analysis tools using Tree-sitter [36], a parsing library for different programming languages. In particular, we construct call graphs by leveraging function names and the numbers of their associated parameters or arguments. The output generated by the indexing agent is stored in a JSON file, enabling incremental analysis as the protocol implementation evolves. This design avoids redundant processing by skipping the analysis of unchanged functions, files, and modules.
To assess the performance of RFCAudit, we conduct experiments to address the following research questions:
  • RQ1: How effective does RFCAudit identify functional bugs in real-world network protocol implementations?
  • RQ2: How does RFCAudit compare to existing approaches?
  • RQ3: What are the runtime and token costs of RFCAudit?
@@ IV-A Dataset
We evaluate RFCAudit on six protocol implementations from two widely-used, open-source network stacks: FRRouting [37] and lwIP [38]. FRRouting (FRR) is an Internet routing protocol suite for Linux and Unix platforms, with over 3.6K stars on GitHub. It is deployed in production by major infrastructure providers such as NVIDIA and Orange, and is distributed through mainstream Linux repositories including Debian and Fedora. lwIP is a lightweight TCP/IP stack designed for resource-constrained environments, with over 1.3K stars on GitHub. It is widely integrated into embedded systems, IoT platforms, and real-time applications, and has been adopted by vendors like Intel, Xilinx, and Freescale. Overall, both FRRouting and lwIP are actively maintained and widely used in practice, making them ideal subjects for assessing RFCAudit across diverse deployment contexts. For each protocol, we collect the corresponding RFC that the implementation is based on as the reference. The protocol implementation sizes range from 5.2 KLoC to 17.3 KLoC, whereas the corresponding RFC documents span between 19 and 54 pages. The detailed information is shown in Table I.
@@
IV-B RQ1: Effectiveness of RFCAudit
IV-B1 The Effectiveness of Bug Detection
@@
To answer RQ1, we apply RFCAudit to each protocol implementation and its corresponding RFC to detect inconsistencies. For every reported issue, we manually verify whether it represents a real violation, and record the counts of true and false positives to compute precision. Since a single underlying issue may result in multiple reported bugs (e.g., when the same property is described across multiple sections of the RFC), we group relevant reports into unique bugs. For each unique bug, we check the project’s issue tracker and pull requests to determine whether it is a previously unreported bug (i.e., new) or a known issue (i.e., old) that has not yet been fixed in the latest version.
Figure 8: Distribution of functional bug categories.
TABLE III: The precision, recall and F1 score of context retrieval. NA indicates that the divisor is 0.
Protocol Properties After Localization After Retrieval #Miss
Function (%) Type (%) Macro (%) Function (%) Type (%) Macro (%)
Babel 15 50.0 / 91.7 / 64.7 NA / 0.0 / NA NA / 0.0 / NA 50.0 / 91.7 / 64.7 100.0 / 100.0 / 100.0 100.0 / 100.0 / 100.0 1
BFD 22 57.1 / 66.7 / 61.5 NA / 0.0 / NA NA / 0.0 / NA 72.7 / 100.0 / 84.2 66.7 / 100.0 / 80.0 100.0 / 100.0 / 100.0 0
NHRP 15 71.4 / 85.7 / 77.9 NA / 0.0 / NA NA / 0.0 / NA 76.9 / 95.2 / 85.1 100.0 / 100.0 / 100.0 100.0 / 100.0 / 100.0 0
RIPng 24 60.7 / 70.8 / 65.4 NA / 0.0 / NA NA / 0.0 / NA 37.3 / 100.0 / 54.3 100.0 / 33.3 / 50.0 NA / 0.0 / NA 0
DHCP 34 60.7 / 68.0 / 64.1 NA / NA / NA NA / NA / NA 65.3 / 94.0 / 77.1 NA / NA / NA NA / NA / NA 0
IGMP 21 52.9 / 75.0 / 62.0 NA / NA / NA NA / 0.0 / / 100.0 / NA / NA / NA 100.0 / 100.0 / 100.0 0
Total 131 59.3 / 71.3 / 65.6 NA / 0.0 / NA NA / 0.0 / NA 56.0 / 95.1 / 70.5 87.5 / 77.8 / 82.4 100.0 / 75.0 / 100.0 1
Results. As shown in Table II, RFCAudit generates a total of 83 bug reports across six protocols, of which 68 are true positives, yielding an overall precision of 81.9%. Grouping related bug reports, we uncovered 47 unique bugs, including 40 new bugs and 7 already reported ones. To date, 20 out of 47 unique bugs have been confirmed or their patches have been approved or merged by the developers. Overall, these results demonstrate that RFCAudit is effective in identifying bugs across diverse protocols. To better understand the diversity of functional bugs detected by RFCAudit, we categorize the 47 unique bugs into seven categories, namely Parsing, Config, Routing, Err Handle, State, Timer, and Demultiplex. As shown in Figure 8, most bugs fall under State (15/47) and Parsing (11/47), reflecting common pitfalls in protocol state management and input validation. RFCAudit also uncovered Timer misconfiguration (e.g., incorrect IGMP delays), Demultiplexing issues (e.g., incorrect BFD session lookup), and Err Handle gaps (e.g., miss error-on-error prevention). This variety highlights the ability of RFCAudit to detect diverse semantic violations specified by the RFCs. Full bug details are available in our artifact [23].
IV-B2 The Effectiveness of Context Retrieval
To assess context retrieval accuracy, we manually construct ground-truth mappings for each RFC property. Each property is annotated with the minimal set of functions, types, and macros required to capture its semantics. Since this annotation is labor-intensive, we sample 30 RFC subsections (5 per protocol) from our datasets, covering a total of 131 properties across six protocols. For each property, we compare the elements retrieved by RFCAudit against the ground truth and measure precision, recall, and F1 score after two stages: (i) localization and (ii) retrieval. We then compute average recall across all sampled properties to assess how effectively RFCAudit collects the necessary context for inconsistency detection. Since both the retrieved element count and ground truth element count can be zero for types and macros, making recall undefined, we use “NA” to indicate these cases. Finally, we examine missed bugs and identify those caused by incomplete retrieval.
Results. Table III reports the context retrieval accuracy across protocols. In the localization stage, RFCAudit retrieves 172 functions, of which 102 are true positives, yielding 59.3% precision and 71.3% recall (102 of 143 ground-truth functions). This stage is limited to function retrieval and, by design, does not include types or macros. After the retrieval stage, precision is 56.0% for functions (136 of 243), 87.5% for types (7 of 8), and 100.0% for macros (7 of 7). More importantly, recall improves significantly—95.1% (136 of 143) for functions, 77.8% (7 of 9) for types, and 75.0% (6 of 8) for macros—with only one bug missed overall, caused by code misinterpretation rather than missing context. In this task, recall is more critical than precision. Low precision only means that unnecessary elements are retrieved, causing additional overhead, while low recall means the needed context is completely missed, preventing the detection of true inconsistencies. Note that achieving high precision is challenging for two reasons: (1) some functions are semantically close to the target property but not truly relevant, making them difficult to filter out, and (2) a single RFC property is often distributed across multiple functions or files, so aggressive filtering risks discarding essential information. For these reasons, our retrieval design explicitly prioritizes recall, even at the cost of including borderline or loosely related items. Precision-oriented refinements are left for future work.
IV-B3 False Positive Analysis
We analyze the 15 false positives reported by RFCAudit to understand their root causes. Based on this analysis, we classify them into four categories: (i) incomplete context retrieval (3 cases), where necessary function, type or macro definitions are not retrieved; (ii) RFC misinterpretation (6 cases), where RFCAudit mistakenly treats optional or incorrect properties as mandatory requirements; (iii) code misinterpretation (5 cases), where RFCAudit retrieves the correct context but fails to correctly interpret the code’s control flow or semantics; and (iv) intentional implementation choice (1 case), where the observed behavior is a deliberate deviation from the RFC. The most common root cause is RFC misinterpretation, including 4 cases where the LLM extracts incorrect properties and 2 cases where the LLM treats optional clauses as mandatory. Such errors may be mitigated by fine-tuning the LLM on RFC texts to improve its ability to extract correct mandatory requirements.
IV-C RQ2: Baseline Comparison
We compare RFCAudit with GitHub Copilot and LTL-Fuzzer [4]. We do not compare with other network protocol bug detection tools, as they either cannot detect functional bugs (e.g., ChatAFL [15], NetLifter [26]), or target narrow scope like parsing (e.g., ParCleanse [30], ParVAL [39]).
IV-C1 Comparison with GitHub Copilot
We compare RFCAudit with GitHub Copilot using three LLMs: Claude 3.7 Sonnet (Thinking Mode), Claude 3.5 Sonnet, and GPT-4o. All are accessed via Copilot’s Workspace Chat interface in VS Code, which supports repository-level queries. To match RFCAudit’s design, we provide Copilot with one RFC section per prompt and ask it to identify deviations from implementations. For Claude 3.7 Sonnet, we manually examine all generated bug reports. However, Claude 3.5 Sonnet and GPT-4o produce hundreds of candidates, making full manual validation impractical. To avoid over-reporting and minimize false positives, we instruct each LLM baseline to only report bugs with 100% confidence. To enable a fair and scalable evaluation, we adopt a sampling-based evaluation. For each protocol, we randomly sample up to twice the number of bug reports identified by RFCAudit (e.g., since RFCAudit generates 16 bug reports for Babel, we sample 32). If the model produces fewer than that, we evaluate all of them. Each sampled case is manually reviewed by a domain expert.
@@ IV-C1 Comparison with GitHub Copilot
Results. Table IV shows the performance of GitHub Copilot under three model configurations. With Claude 3.7 Sonnet (no sampling), it generated 215 bug reports, of which 78 were true positives (36.3% precision), uncovering 37 unique bugs (29 new, 8 old). Claude 3.5 Sonnet achieved 19.9% precision over 538 inconsistencies and found 26 bugs. GPT-4o returned the most bug reports (1,315) but the lowest precision (11.4%) and the fewest bugs (16). Overall, all three baselines suffer from low precision and high validation overhead. In contrast, RFCAudit achieves 81.9% precision and discovers 47 bugs in total (40 new) with far less manual effort (Table II). RFCAudit consistently reports the lowest false-positive rates and the highest bug counts across all protocols. Notably, RFCAudit ’s use of Claude 3.5 (a non-reasoning model) still outperforms Claude 3.7 (a reasoning model).
IV-C2 Comparison with LTL-Fuzzer
We compare RFCAudit with LTL-Fuzzer [4], which checks protocol implementations against LTL properties via instrumentation and fuzzing. Setting up LTL-Fuzzer requires substantial manual efforts: translating informal specifications into LTL formulas and mapping each atomic predicate to code locations. Consequently, it mainly captures event-ordering temporal properties but cannot naturally express data-format checks, arithmetic constraints, or timing requirements. We applied RFCAudit to the 15 zero-day bugs reported in LTL-Fuzzer’s evaluation and also examined whether the 47 violated properties from our dataset could be encoded for LTL-Fuzzer properties.
Results. RFCAudit successfully detected 12 of the 15 zero-days originally reported by LTL-Fuzzer, without requiring manual property construction or predicate-to-location mapping. These included 7 in TinyDTLS and 5 in Contiki-Telnet. In contrast, when we attempted to encode the 47 violated properties from our dataset as LTL-Fuzzer properties, only 11 could be expressed; the remaining 36 required constraints beyond plain LTL or lacked precise predicate locations. These results indicate that RFCAudit is effective in automatically detecting temporal property violations while extending to a broader range of functional inconsistencies.
IV-D RQ3: Efficiency of RFCAudit